How do you solve slope intercept form
2×+y=22

Answer by macston(5194)

.
Slope intercept form: y=mx+b
.
Move every thing but y to the right side.
.
2x+y=22 . Subtract 2x from each side.
y=22-2x . Shuffle the right side to get term containing x first.
.
y=-2x+22
.